OpenStudy (anonymous):

Estimate the sum or difference. 15 2/15 - 15 1/20 A 2 B 1 C 0

OpenStudy (wolf1728):

Are you sure of those answers? The total should be about 30

TheSmartOne (thesmartone):

^^

OpenStudy (anonymous):

yes I am sure

OpenStudy (wolf1728):

Then I guess you must estimate the diffference (that "+" in the equation fooled me)

OpenStudy (anonymous):

hold on sorry

OpenStudy (wolf1728):

Estimate 15 2/15 - 15 1/20

TheSmartOne (thesmartone):

That is probably what it is ^^

OpenStudy (wolf1728):

Well the 15's cancel out and you must now solve 2/15 - 1/20 which easily rounds to zero

OpenStudy (wolf1728):

You could get an exact answer without a clculator: 2/15 = 8/60 1/20 = 3/60 8/60 -3/60 = 5/60 5/60 = 1/12